OpenStudy (anonymous):

Please Help !!! Identify the oblique asymptote of f(x) = quantity 2 x squared plus 3 x plus 8 over quantity x plus 3 y = 2x - 3 y = 2x + 9 y = 0 No oblique asymptote

OpenStudy (anonymous):

@helpme1.2

OpenStudy (anonymous):

@abb0t

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

@beccaboo333

OpenStudy (anonymous):

@Loser66

OpenStudy (beccaboo333):

Sorry I can't really help you here. Try posting the link to this in the chats?

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

you will get a quotient and a remainder ignore the remainder, the quotient is the slant asympote

OpenStudy (anonymous):

huh ? i don't get it, i'm completely lost. @satellite73

OpenStudy (anonymous):

you know what i mean by "divide" we can do it if you like

OpenStudy (anonymous):

ok can you walk me through

OpenStudy (anonymous):

\[\frac{2x^2+3x+3}{x+8}\] right?

OpenStudy (anonymous):

it is hard to write division here do you know how to do synthetic division? that is much easier

OpenStudy (anonymous):

not quite, its 2x^2 + 3x+8 over x + 3

OpenStudy (anonymous):

\[\frac{2x^2+3x+8}{x+3}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you know synthetic division?

OpenStudy (anonymous):

no, I''m not very good at math. and they make us use calculators in schools here

OpenStudy (anonymous):

2 3 8 -3 ___________

OpenStudy (anonymous):

bring down the 2 2 3 8 -3 _____________ 2

OpenStudy (anonymous):

then \(-3\times 2=-6) 2 3 8 -3 -6 _____________ 2

OpenStudy (anonymous):

add 2 3 8 -3 -6 _____________ 2 -3

OpenStudy (anonymous):

we can actually stop right there the bottom line 2 , -3 means the quotient is \(2x-3\) and your slant asymptote is \[y=2x-3\]

OpenStudy (anonymous):

there is a remainder, but you can ignore it for the slant asymptote if you don't do it this way you have to do long division x + 3 |2x^2 + 3 +8 etc can't really write it in this format

OpenStudy (anonymous):

hmm, so it's A ?

OpenStudy (anonymous):

i have no idea what A is but i am pretty sure it is \(y=2x-3\)

OpenStudy (anonymous):

okay , thanks so much

OpenStudy (anonymous):

yw